the question i am having problem with is: 5x-y=10, what is the value of x. can you please help me with this problem.
Answer by serenebarril(1)

5x-y=10
5x-y+y=10+y - Addition property of equality
5x/5=(10+y)/5-Division Property of Equality
x=(10+y)/5
note: you can't get an exact value for x because you need at least two equations when you have two variables